# Representations of categories of finite relational structures and associated endomorphism monoids

### 18 Mar 2026 v1

@@ Abstract.
We develop a unified representation theory for the categories of finite substructures of highly homogeneous relational structures classified by Cameron. For any commutative coefficient ring $k$, we extend the classical Dold–Kan correspondence to this setting—with the sole exception of the category $\mathrm{FA}$}—and prove that finitely generated representations are noetherian (resp., artinian) when $k$ is noetherian (resp., artinian).
When $k$ is a field of characteristic 0, we obtain a precise structural description of these representation categories. We classify irreducible representations, show that every indecomposable standard modules either is irreducible (the regular case) or has length 2 (the singular case), and establish a (direct sum or triangular) decomposition into a singular component governed by classical Dold–Kan theory and a regular component exhibiting semisimplicity or representation stability.
Finally, we establish a monoidal generalization of Artin’s reconstruction theorem for topological groups, proving an equivalence between uniformly continuous representations of infinite endomorphism monoids and sheaves on the associated finite substructure categories. In the special case where the endomorphism monoid is a permutation group, our result recovers Artin’s theorem.
@@ 1.1. Motivation
In a different direction, the representation theory of categories of finite sets with structured injective morphisms has developed rapidly in recent years, particularly through the theory of $\mathrm{FI}$}-modules introduced by Church–Ellenberg–Farb [5] and further developed by many authors; see for instances [6, 17, 18, 19, 20, 22, 31, 37, 38, 43, 45, 48] and the surveys [46, 47]. These works establish deep structural results, including noetherianity, local self-injectivity, and connections to representation stability, polynomial functors and functor homology [11, 12]. By contrast, categories in which injective and surjective morphisms coexist symmetrically remain largely outside the scope of current structural theories. Notable exceptions include the category $\mathrm{FA}$} of finite sets and maps and the category ${\mathrm{VA}}_{q}$ of finite dimension vector spaces over a finite field $\mathbb{F}_{q}$ and linear maps. Their representation theory has been investigated by Wiltshire-Gordon, Powell–Vespa, and Kuhn [27, 41, 42, 51].
Our aim is to develop a representation theory for combinatorial categories in which injective and surjective morphisms are treated on equal footing. A natural source of such categories arises from finite substructure categories associated with oligomorphic group actions [4]. On one hand, these categories include many important examples previously studied in [1, 15]; on the other hand, this framework provides a structural bridge between representations of small categories and continuous representations of their associated automorphism groups or endomorphism monoids, in the sense of categorical reconstruction [10, 30].
As a first step, we focus on the class of highly homogeneous relational structures, classified by Cameron [3, 4]. Cameron showed that there exist exactly five countable relational structures whose automorphism groups act highly homogeneously. Each such structure yields a skeletal category of finite substructures and structure-preserving maps, giving rise to the five categories
@@
The purpose of this paper is to initiate a systematic study of representations of these five categories, addressing several conceptual gaps. First, existing results are largely fragmentary and rely on case-by-case combinatorial arguments, with no uniform framework treating all five categories simultaneously. Second, the appearance of simplicial phenomena—such as the Dold–Kan correspondence—has not been explained intrinsically from the ambient categorical structure. Third, the relationship between representations of these combinatorial categories and representations of associated infinite endomorphism monoids has not been clarified at a structural level.
@@ 1.2. Notations
Fix a commutative ring $k$. A representation of $\mathscr{C}$} (or a $\mathscr{C}$}-module) $V$ is a covariant functor from $\mathscr{C}$} to $k{\textrm{-}\mathrm{Mod}}$, the category of $k$-modules. Denote by $P_{n}$ the representable functor $k{\mathscr{C}}(n,-)$. Following [9], we define the standard module $\Delta_{n}$ as the quotient of $P_{n}$ by the submodule spanned by all non-injective morphisms with source $[n]$. The generalized Reedy structure ensures that $P_{n}$ admits a finite filtration with standard modules as its factors. In the case where $k$ is a field of characteristic 0 and $\lambda$ is an irreducible $kG_{n}$-module, we define the indecomposable standard module
$\Delta_{n,\lambda}=\Delta_{n}\otimes_{kG_{n}}\lambda,$
which is a direct summand of $\Delta_{n}$.
@@ 1.3. Main results
While this result is classical for the simplex category and known for the cyclic category [14, Corollary 6.6], to our knowledge this unified treatment is new for the categories $\mathrm{BA}$} and $\mathrm{SA}$}. In contrast to the result by Kaygun–Kaya [26], which applies normalization procedures to crossed simplicial groups within a model-categorical context, our work focuses on the representation theory of these categories.
@@
This result is well known for $\mathrm{OA}$} and can be deduced for $\mathrm{CA}$} from [14, Corollary 6.6]. For $\mathrm{FA}$}, the noetherian fact is established in [48, Corollary 7.3.5], and the finite length result has been established in [42, Corollary 4.33] and [51, Theorem 5.3] under the assumption that $k$ is a field.
@@
When $k$ is a field of characteristic 0, the structure becomes more refined. In this case, each standard module admits a canonical decomposition
$\Delta_{n}\cong\bigoplus_{\lambda\in\operatorname{Irr}(G_{n})}(\Delta_{n,\lambda})^{\oplus\dim\lambda},$
where $\operatorname{Irr}(G_{n})$ is a complete set of non-isomorphic irreducible $kG_{n}$-modules. We call $\lambda\in\operatorname{Irr}(G_{n})$ singular if it appears in the classification of Proposition 5.4, that is, there exists a nonzero homomorphism from $\Delta_{n+1}$ to $\Delta_{n,\lambda}$. Otherwise, $\lambda$ is called regular. More explicitly, $\lambda\in\operatorname{Irr}(G_{n})$ is singular if and only if it is isomorphic to the one-dimensional space spanned by the alternating sum of the vertices in a connected component of the mutation graph $\Gamma_{n-1,n}$. These simple combinatorial observations yield a sharp structural dichotomy.
Theorem 1.7 (Propositions 5.5 and 5.6).
Let $k$ be a field of characteristic 0. Then $\Delta_{n,\lambda}$ is irreducible if $\lambda$ is regular, and has length $2$ while $\lambda$ is singular.
@@
Let $k$ be a field of characteristic 0. Isomorphism classes of irreducible $\mathscr{C}$}-modules are parameterized by pairs $(n,\lambda)$ with $n\in\mathbb{Z}_{+}$ and $\lambda\in\mathrm{Irr}(G_{n})$. If $\lambda$ is regular, the corresponding irreducible $\mathscr{C}$}-module is $\Delta_{n,\lambda}$; if $\lambda$ is singular, it is the top of $\Delta_{n,\lambda}$.
@@
For the category $\mathrm{OA}$}, this classification of irreducible modules holds over any field and is considered classical. For $\mathrm{BA}$}, the result remains valid over any field of characteristic different from $2$. For the category $\mathrm{FA}$}, a complete classification of irreducible representations has been previously established using different methods; see [42, Theorem 1.4].
Using the regular–singular dichotomy, we obtain a refined form of the Dold–Kan correspondence. More precisely, we construct a $k$-linear category $\mathscr{K}$ whose representation category is equivalent to ${\mathscr{C}}{\textrm{-}\mathrm{Mod}}$, but whose structure is substantially simpler than the category appearing in Theorem 1.1. For $\mathrm{BA}$}, the category $\mathscr{K}$ is essentially the direct sum of two copies of the $k$-linear category arising in the classical Dold–Kan correspondence for $\mathrm{OA}$}.
For the remaining categories, the structure exhibits new and genuinely different behavior. Specifically, one has:
  • If ${\mathscr{C}}={\mathrm{CA}}$ or $\mathrm{SA}$}, then every regular indecomposable standard module is irreducible and projective. As a consequence, we have a decomposition
@@
Our final main result places the representation theory of $\mathscr{C}$} into a genuinely infinite-dimensional and structural framework. Explicitly, let $\mathscr{E}\leqslant\operatorname{End}(\Omega,\mathcal{R}_{G})$ be an endomorphism monoid of the relational structure $(\Omega,\mathcal{R}_{G})$ induced by the action (which might not be highly homogeneous) of a group $G$ on an infinite set $\Omega$. We equip $\mathscr{E}$ with the topology of pointwise convergence. A $k\mathscr{E}$-module $M$ is said to be continuous if the action of $\mathscr{E}$ on $M$ is continuous with respect to the discrete topology on $M$; it is uniformly continuous if the action of $\mathscr{E}$ is uniformly continuous with respect to the natural uniform structure on $\mathscr{E}$ induced by pointwise convergence.
Let $\mathscr{C}$} be the category of finite substructures of $(\Omega,\mathcal{R}_{G})$ whose morphisms are restrictions of endomorphisms in $\mathscr{E}$. A $\mathscr{C}$}-module $V$ is said to be torsion if for every $A\in\operatorname{Ob}({\mathscr{C}})$ and every $v\in V(A)$, there is an inclusion $\iota_{A,B}:A\hookrightarrow B$ such that $\iota_{A,B}\cdot v=0$. We also define the Grothendieck topology $J$ on ${\mathscr{C}}^{{\mathrm{op}}}$ as follows: for every $A\in\operatorname{Ob}({\mathscr{C}})$, a sieve is a covering sieve in $J(A)$ if it contains an inclusion morphism with source $A$.
@@
This equivalence provides a systematic framework connecting three different perspectives: representations of combinatorial categories, uniformly continuous representations of infinite endomorphism monoids, and their interpretation via sheaf-theoretic localization. In this sense, it may be viewed as a monoidal analogue of Artin’s classical reconstruction theorem for topological groups (see [34, III.9]), extending the reconstruction principle from invertible symmetries to general endomorphism monoids. The equivalence therefore allows one to translate structural and representation-theoretic questions about infinite transformation monoids into problems about representations of categories of finite structures, where powerful algebraic techniques are available. We discuss this relationship in more detail in Subsection 7.4.
@@
where the colimit is taken over the poset of finite substructures ordered by inclusion. Under this assumption, no nonzero $\mathscr{C}$}-module is torsion with respect to inclusions, so the localization step disappears and the correspondence becomes completely canonical. This condition is satisfied by the five combinatorial categories arising from highly homogeneous relational structures. Consequently, the theorem allows one to apply the rich theory of representations of these combinatorial categories to the study of uniformly continuous representations of the corresponding infinite endomorphism monoids.
A striking feature of the theory is that very different monoids have identical uniformly continuous representation theory. Explicitly, for any homogeneous relational structure, the automorphism group shares the same uniformly continuous representation theory as the monoid of injective endomorphisms; in this case, every continuous representation of the monoid is automatically uniformly continuous. For highly homogeneous structures, the full endomorphism monoid and the submonoid of surjective endomorphisms still share the same uniformly continuous representations, although continuous representations of these monoids may not be uniformly continuous. These examples illustrate that, despite the combinatorial differences, their uniformly continuous representation theory can be indistinguishable. For details, see Subsection 7.5.
@@ 1.4. Organization and conventions
The paper is organized as follows. In Section 2, we study categories of finite substructures arising from group actions on sets and prove that they admit generalized Reedy structures in full generality. Section 3 introduces mutation graphs, a key combinatorial tool used to compute Hom- and Ext-spaces between standard modules. In Section 4, we establish the generalized Dold–Kan correspondence and the equivalence with the normalized category $\mathscr{K}$ for all Cameron categories except $\mathrm{FA}$}. Section 5 is devoted to explicit computations of Hom-spaces between standard modules and a classification of irreducible modules. In Section 6, we present further representation-theoretic results for $\mathrm{CA}$}, $\mathrm{SA}$}, and $\mathrm{FA}$}, including the refined Dold–Kan correspondence. Finally, in Section 7, we investigate the relationship between uniformly continuous representations of endomorphism monoids and representations of the associated categories of finite substructures, establishing a monoidal extension of Artin’s theorem.
@@ 2. Reedy Structure
Throughout this paper, let $\Omega$ be an infinite set, and let $G\times\Omega\to\Omega$ be an action of a group $G$ on $\Omega$. This action induces a canonical relational structure $(\Omega,\mathcal{R}_{G})$: for each $n\geqslant 1$, two $n$-tuples in $\Omega^{n}$ are equivalent if and only if they lie in the same $G$-orbit. This structure is homogeneous in the following sense: if $A$ and $B$ are finite subsets of $\Omega$, equipped with the induced relational structure, and if $\sigma:A\to B$ is an isomorphism of relational structures, then $\sigma$ extends to an element of GG. [^1] More precisely, $\sigma$ extends to an element of $\operatorname{Aut}(\Omega,\mathcal{R}_{G})$. Since $G$ is dense in $\operatorname{Aut}(\Omega,\mathcal{R}_{G})$ with respect to the permutation topology, one may choose such an extension inside $G$. See, for example, [2] for details. In fact, every embedding $A\hookrightarrow B$ of finite substructures extends to an element of $G$.
Let $\mathscr{C}$} be a skeletal full subcategory of the category of finite substructures of $\Omega$ and relation-preserving maps. Thus every finite substructure of $\Omega$ is isomorphic to a unique object in $\mathscr{C}$}. Denote by $\mathscr{C}$$^{+}$ the wide subcategory consisting of injective morphisms, and by $\mathscr{C}$$^{-}$ the wide subcategory consisting of surjective morphisms. We will show that $\mathscr{C}$} is a generalized Reedy category in the sense of Berger–Moerdijk [1]. For an object $A\in{\mathscr{C}}$, denote by $G_{A}$ its automorphism group; equivalently,
@@
Firstly, we show the existence of such a factorization. Define an equivalence relation $\sim$ on the underlying set of $B$ by
@@
Since $f$ is a morphism of relational structures, the relation $\sim$ is compatible with all relations on $B$, so the quotient set $C=B/\sim$ inherits a natural relational structure. Without loss of generality we can assume that $C$ is an object in $\mathscr{C}$}. Thus the quotient map $f^{-}:B\twoheadrightarrow C$ is a surjective morphism in $\mathscr{C}$}. The universal property of the quotient now yields a unique map $f^{+}:C\to A$ such that $f=f^{+}\circ f^{-}$. Since distinct equivalence classes map to distinct points of $A$, the map $f^{+}$ is injective. It is straightforward to verify that $f^{+}$ preserves the relational structure, hence is an injective morphism in $\mathscr{C}$}. This gives the required factorization.
Now we show the uniqueness of the factorization. Suppose $f=g^{+}\circ g^{-}$ is another factorization with $g^{-}$ surjective and $g^{+}$ injective. Then for $x,y\in B$, one has
$g^{-}(x)=g^{-}(y)\;\Longrightarrow\;f(x)=f(y),$
so $\ker(g^{-})=\sim$. By the universal property of quotient structures, there exists a unique isomorphism $\sigma:C\to C$ such that $g^{-}=\sigma\circ f^{-}$. Substituting into $f=g^{+}\circ g^{-}$ and using surjectivity of $f^{-}$, we obtain $f^{+}=g^{+}\circ\sigma$. Thus $\sigma\in G_{C}$ makes the diagram commute. The uniqueness of $\sigma$ follows from Lemma 2.1, since the right action of $G_{C}$ on ${\mathscr{C}}^{-}(B,C)$ is free. ∎
@@
We collect some fundamental properties of these modules in the following proposition, which follow from Lemma 2.1, Proposition 2.3, and [9, Lemmas 2.10 and 3.4, and Theorem 3.6]. Set ${\mathscr{C}}^{0}={\mathscr{C}}^{+}\cap{\mathscr{C}}^{-}$, the wide subcategory of $\mathscr{C}$} consisting of isomorphisms.

### 25 Apr 2026 v3

@@ Abstract.
We develop a unified representation theory for the categories of finite subsets and relation-preserving maps of highly homogeneous relational structures classified by Cameron. For any commutative coefficient ring $k$, we extend the classical Dold–Kan correspondence to this setting—with the sole exception of the category $\mathrm{FA}$}—and prove that finitely generated representations are noetherian (resp., artinian) when $k$ is noetherian (resp., artinian).
When $k$ is a field, we obtain a precise structural description of these representation categories. We classify irreducible representations, showing that canonical quotients of indecomposable standard modules are either irreducible (the regular case) or has length 2 (the singular case). In the case that $k$ has characteristic 0, we establish a (direct sum or triangular) decomposition into a singular component governed by classical Dold–Kan theory and a regular component exhibiting semisimplicity or representation stability.
Finally, we establish a monoidal generalization of Artin’s reconstruction theorem for topological groups, proving an equivalence between uniformly continuous representations of infinite topological transformation monoids and sheaves on the associated categories of finite subsets. In the special case where the transformation monoid is a permutation group, our result recovers Artin’s theorem.
@@ 1.1. Motivation
In a different direction, the representation theory of categories of finite sets with structured injective morphisms has developed rapidly in recent years, particularly through the theory of $\mathrm{FI}$}-modules introduced by Church–Ellenberg–Farb [6] and further developed by many authors; see for instances [7, 19, 20, 21, 22, 24, 33, 39, 40, 45, 48, 51] and the surveys [49, 50]. These works establish deep structural results, including noetherianity, local self-injectivity, and connections to representation stability, polynomial functors and functor homology [13, 14]. By contrast, categories in which injective and surjective morphisms coexist symmetrically remain largely outside the scope of current structural theories. Notable exceptions include the category $\mathrm{FA}$} of finite sets and maps and the category ${\mathrm{VA}}_{q}$ of finite dimension vector spaces over a finite field $\mathbb{F}_{q}$ and linear maps. Their representation theory has been investigated by Kuhn, Powell, Vespa, and Wiltshire-Gordon [29, 43, 44, 54]. Moreover, representation theory of $\mathrm{FA}$} has been applied to study holomorphic forms on moduli spaces of stable curves by Canning–Larson–Payne–Willwacher [5].
Our aim is to develop a representation theory for combinatorial categories in which injective and surjective morphisms are treated on equal footing. A natural source of such categories arises from categories of finite subsets and relation-preserving maps associated with oligomorphic group actions [4]. On one hand, these categories include many important examples previously studied in [1, 17]; on the other hand, this framework provides a structural bridge between representations of small categories and continuous representations of their associated automorphism groups or endomorphism monoids, in the sense of categorical reconstruction [12, 32].
As a first step, we focus on the class of highly homogeneous relational structures, classified by Cameron [3, 4]. Cameron showed that there exist exactly five highly homogeneous relations on a countable set. Each such relation yields a skeletal category of finite subsets and relation-preserving maps, giving rise to the five categories
@@
The purpose of this paper is to initiate a systematic study of representations of these five categories, addressing several conceptual gaps. First, existing results are largely fragmentary and rely on case-by-case combinatorial arguments, with no uniform framework treating all five categories simultaneously. Second, the appearance of simplicial phenomena—such as the Dold–Kan correspondence—has not been explained intrinsically from the ambient categorical structure. Third, the relationship between representations of these combinatorial categories and representations of associated infinite transformation monoids has not been clarified at a structural level.
@@ 1.2. Notations
Fix a commutative ring $k$. A representation of $\mathscr{C}$} (or a $\mathscr{C}$}-module) $V$ is a covariant functor from $\mathscr{C}$} to $k{\textrm{-}\mathrm{Mod}}$, the category of $k$-modules. Denote by $P_{n}$ the representable functor $k{\mathscr{C}}(n,-)$. We define the standard module $\Delta_{n}$ as the quotient of $P_{n}$ by the submodule spanned by all non-injective morphisms with source $[n]$, an analogue of the classical notion in algebraic representation theory: Dalezois-Št’ovíček [10] define $k$-linear Reedy categories with $k$ a field, which can be viewed infinite quasi-hereditary algebras; inspired by their work, Di-Li-Liang [11] define generalized $k$-linear Reedy categories for arbitrary commutative rings $k$, which can be viewed as infinite standardly stratified algebras. The generalized Reedy structure ensures that $P_{n}$ admits a finite filtration with standard modules as its factors. In the case where $k$ is a field and $\lambda$ is an irreducible $kG_{n}$-module, we define the indecomposable standard module
$\Delta_{n,\lambda}=\Delta_{n}\otimes_{kG_{n}}kG_{n}e_{\lambda},$
where $e_{\lambda}\in kG_{n}$ is the primitive idempotent such that $kG_{n}e_{\lambda}$ is the projective cover of $\lambda$. Correspondingly, we define
$\overline{\Delta}_{n,\lambda}=\Delta_{n}\otimes_{kG_{n}}\lambda,$
which is a quotient of $\Delta_{n,\lambda}$, and is isomorphic to $\Delta_{n,\lambda}$ when $k$ has characteristic 0.
@@ 1.3. Main results
While this result is classical for the simplex category and known for the cyclic category [16, Corollary 6.6], to our knowledge this unified treatment is new for the categories $\mathrm{BA}$} and $\mathrm{SA}$}. In contrast to the result by Kaygun–Kaya [28], which applies normalization procedures to crossed simplicial groups within a model-categorical context, our work focuses on the representation theory of these categories.
@@
This result is well known for $\mathrm{OA}$} and can be deduced for $\mathrm{CA}$} from [16, Corollary 6.6]. For $\mathrm{FA}$}, the noetherian fact is established in [51, Corollary 7.3.5], and the finite length result has been established in [44, Corollary 4.33] and [54, Theorem 5.3] under the assumption that $k$ is a field.
@@
When $k$ is a field, the structure becomes more refined. In this case, each standard module admits a canonical decomposition
$\Delta_{n}\cong\bigoplus_{\lambda\in\operatorname{Irr}(G_{n})}(\Delta_{n,\lambda})^{\oplus a_{\lambda}},$
where $\operatorname{Irr}(G_{n})$ is a complete set of non-isomorphic irreducible $kG_{n}$-modules. We call $\lambda$ singular if it appears in the classification of Proposition 5.4, that is, there exists a nonzero homomorphism from $\Delta_{n+1}$ to $\Delta_{n,\lambda}$. Otherwise, $\lambda$ is called regular. More explicitly, $\lambda\in\operatorname{Irr}(G_{n})$ is singular if and only if it is isomorphic to the one-dimensional space spanned by the alternating sum of the vertices in a connected component of the mutation graph $\Gamma_{n-1,n}$.
These simple combinatorial observations yield a sharp structural dichotomy.
Theorem 1.7 (Propositions 5.6 and 5.9).
Let $k$ be a field. Then $\overline{\Delta}_{n,\lambda}$ is irreducible if $\lambda$ is regular. If $\lambda$ is singular, then $\overline{\Delta}_{n,\lambda}$ has length $2$.
@@
Let $k$ be a field. Isomorphism classes of irreducible $\mathscr{C}$}-modules are parameterized by pairs $(n,\lambda)$ with $n\in\mathbb{Z}_{+}$ and $\lambda\in\mathrm{Irr}(G_{n})$. If $\lambda$ is regular, the corresponding irreducible $\mathscr{C}$}-module is $\overline{\Delta}_{n,\lambda}$; if $\lambda$ is singular, it is the top of $\overline{\Delta}_{n,\lambda}$.
@@
For the category $\mathrm{OA}$}, this classification of irreducible modules is classical. For $\mathrm{FA}$}, the above classification has been established via different methods under the assumption that $k$ has characteristic 0; see [44, Theorem 1.4], [47, Theorem 4.1], or [54, Theorem 5.5]. Our approach extends these classifications to fields of arbitrary characteristic.
Using the regular–singular dichotomy, we obtain a refined form of the Dold–Kan correspondence in the case that $k$ is a field of characteristic 0. More precisely, we construct a $k$-linear category $\mathscr{K}$ whose representation category is equivalent to ${\mathscr{C}}{\textrm{-}\mathrm{Mod}}$, but whose structure is substantially simpler than the category appearing in Theorem 1.1. For $\mathrm{BA}$}, the category $\mathscr{K}$ is essentially the direct sum of two copies of the $k$-linear category arising in the classical Dold–Kan correspondence for $\mathrm{OA}$}. For the remaining categories, the structure exhibits new and genuinely different behavior. Specifically, one has:
  • If ${\mathscr{C}}={\mathrm{CA}}$ or $\mathrm{SA}$}, then every regular indecomposable standard module is projective (this is true for any field; see Proposition 6.3) and irreducible. As a consequence, we have a decomposition
@@
Our final main result places the representation theory of $\mathscr{C}$} into a genuinely infinite-dimensional and structural framework. Explicitly, let $\mathscr{E}\leqslant\Omega^{\Omega}$ be a transformation monoid. We equip $\mathscr{E}$ with the topology of pointwise convergence. A $k\mathscr{E}$-module $M$ is said to be continuous if the action of $\mathscr{E}$ on $M$ is continuous with respect to the discrete topology on $M$; it is uniformly continuous if the action of $\mathscr{E}$ is uniformly continuous with respect to the natural uniform structure on $\mathscr{E}$ induced by pointwise convergence.
Let $\mathscr{C}$} be the category of finite subsets of $\Omega$ whose morphisms are restrictions of elements in $\mathscr{E}$. A $\mathscr{C}$}-module $V$ is said to be torsion if for every $A\in\operatorname{Ob}({\mathscr{C}})$ and every $v\in V(A)$, there is an inclusion $\iota_{A,B}:A\hookrightarrow B$ such that $\iota_{A,B}\cdot v=0$. We also define the Grothendieck topology $J$ on ${\mathscr{C}}^{{\mathrm{op}}}$ as follows: for every $A\in\operatorname{Ob}({\mathscr{C}})$, a sieve is a covering sieve in $J(A)$ if it contains an inclusion morphism with source $A$.
@@
This equivalence provides a systematic framework connecting three different perspectives: representations of combinatorial categories, uniformly continuous representations of infinite transformation monoids, and their interpretation via sheaf-theoretic localization. In this sense, it may be viewed as a monoidal analogue of Artin’s classical reconstruction theorem for topological groups (see [36, III.9]), extending the reconstruction principle from invertible symmetries to general transformation monoids. The equivalence therefore allows one to translate structural and representation-theoretic questions about infinite transformation monoids into problems about representations of categories $\mathscr{C}$} of finite subsets of $\Omega$, where powerful algebraic techniques are available. We discuss this relationship in more detail in Subsection 7.4.
@@
where the colimit is taken over the poset of finite subsets ordered by inclusion. Under this assumption, no nonzero $\mathscr{C}$}-module is torsion with respect to inclusions, so the localization step disappears and the correspondence becomes completely canonical. This condition is satisfied by the five combinatorial categories arising from highly homogeneous relations (when $\mathscr{E}$ is taken to be the monoid of maps preserving the underlying defining relations on $\Omega$). Consequently, the theorem allows one to apply the rich theory of representations of these combinatorial categories to the study of uniformly continuous representations of the corresponding infinite transformation monoids.
A striking feature of the theory is that very different monoids can have identical uniformly continuous representation theory. Explicitly, for any highly homogeneous relation on $\Omega$, the automorphism group shares the same uniformly continuous representation theory as the monoid of injective relation-preserving maps; in this case, every continuous representation of the monoid is automatically uniformly continuous. Similarly, the full transformation monoid consisting of all relation-preserving maps and the submonoid of surjective such maps still share the same uniformly continuous representations, although continuous representations of these monoids may fail to be uniformly continuous. These examples illustrate that, despite substantial combinatorial differences, their uniformly continuous representation theory can be indistinguishable. For details, see Subsection 7.5.
@@ 1.4. Organization and conventions
The paper is organized as follows. In Section 2, we study categories of finite subsets and relation-preserving maps and prove that they admit generalized Reedy structures in full generality. Section 3 introduces mutation graphs, a key combinatorial tool used to compute Hom- and Ext-spaces between standard modules. In Section 4, we establish the generalized Dold–Kan correspondence and the equivalence with the normalized category $\mathscr{K}$ for all Cameron categories except $\mathrm{FA}$}. Section 5 is devoted to explicit computations of Hom-spaces between standard modules and a classification of irreducible modules. In Section 6, we present further representation-theoretic results for $\mathrm{CA}$}, $\mathrm{SA}$}, and $\mathrm{FA}$}, including the refined Dold–Kan correspondence. Finally, in Section 7, we investigate the relationship between uniformly continuous representations of endomorphism monoids and representations of the associated categories of finite subsets, establishing a monoidal extension of Artin’s theorem.
@@ 2. Reedy Structure
Throughout this paper, let $\Omega$ be an infinite set equipped with an underlying relation $\mathcal{R}_{\Omega}$. Let $\mathscr{C}$} be a skeletal full subcategory of the category whose objects are finite subsets of $\Omega$ equipped with relations inherited from $\mathcal{R}_{\Omega}$, and whose morphisms are relation-preserving maps. Thus every finite subset of $\Omega$ is isomorphic to a unique object in $\mathscr{C}$}. Denote by $\mathscr{C}$$^{+}$ the wide subcategory consisting of injective morphisms, and by $\mathscr{C}$$^{-}$ the wide subcategory consisting of surjective morphisms.
We will show that $\mathscr{C}$} is a generalized Reedy category in the sense of Berger–Moerdijk [1]. For an object $A\in{\mathscr{C}}$, denote by $G_{A}$ its automorphism group; equivalently,
@@
We first show the existence of the factorization. Define an equivalence relation $\sim$ on the underlying set of $B$ by
@@
Let $C=B/\sim$ be the set of equivalence classes. We equip $C$ with a relational structure as follows: for any relation symbol $R$ and any $x_{1}],\dots,[x_{n}]\in C$, define
$R_{C}([x_{1}],\dots,[x_{n}])\quad\Longleftrightarrow\quad R_{A}(f(x_{1}),\dots,f(x_{n})).$
This is well-defined, since if $x_{i}\sim x_{i}^{\prime}$ for all $i$, then $f(x_{i})=f(x_{i}^{\prime})$, so the right-hand side does not depend on the choice of representatives.
Without loss of generality, we may assume that $C$ is an object of $\mathscr{C}$}. Let $f^{-}:B\twoheadrightarrow C$ be the canonical projection. By construction, $f^{-}$ is surjective. Moreover, if $R_{B}(x_{1},\dots,x_{n})$ holds in $B$, then since $f$ is relation-preserving, $R_{A}(f(x_{1}),\dots,f(x_{n}))$ holds in $A$, and hence $R_{C}([x_{1}],\dots,[x_{n}])$ holds in $C$. Thus $f^{-}$ is relation-preserving, so $f^{-}\in{\mathscr{C}}^{-}$.
Next, define $f^{+}:C\to A$ by $f^{+}([x])=f(x)$. This is well-defined by definition of $\sim$, and by construction we have $f=f^{+}\circ f^{-}$. It is clear that $f^{+}$ is injective. Furthermore, if $R_{C}([x_{1}],\dots,[x_{n}])$ holds, then by definition $R_{A}(f(x_{1}),\dots,f(x_{n}))$ holds, so $R_{A}(f^{+}([x_{1}]),\dots,f^{+}([x_{n}]))$ holds as well. Thus $f^{+}$ is relation-preserving, so $f^{+}\in{\mathscr{C}}^{+}$. This gives the required factorization.
For uniqueness, suppose $f=g^{+}\circ g^{-}$ is another factorization. Then for $x,y\in B$,
$g^{-}(x)=g^{-}(y)\quad\Longrightarrow\quad f(x)=f(y),$
so $\ker(g^{-})=\sim$. By the universal property, there exists a unique map $\sigma:C\to C$ such that
$g^{-}=\sigma\circ f^{-}.$
Since $g^{-}=\sigma\circ f^{-}$ and $f^{-}$ is surjective, it follows that $\sigma$ is relation-preserving and surjective; from
$f=f^{+}\circ f^{-}=g^{+}\circ g^{-}=g^{+}\circ\sigma\circ f^{-}$
and the surjectivity of $f^{-}$, we deduce that $f^{+}=g^{+}\circ\sigma$ and the injectivity of $\sigma$. Thus $\sigma$ is contained in $G_{C}$ and makes the diagram commute. ∎
@@
We collect some fundamental properties of these modules in the following proposition, which follow from Lemma 2.1, Proposition 2.3, and [11, Lemmas 2.10 and 3.4, and Theorem 3.6]. Set ${\mathscr{C}}^{0}={\mathscr{C}}^{+}\cap{\mathscr{C}}^{-}$, the wide subcategory of $\mathscr{C}$} consisting of isomorphisms.